What is 5y-6x when y equals 9 and x equals 4?
Answer by oberobic(2304)

Just substitute the values and multiply
.
5y - 6x = ?
.
y = 9 (given)
x = 4 (given)
.
5*9 - 6*4 = 45 -24
45 -24 = 21
.
5y -6x = 21, when y=9 and x=4
.
Done.